# Multiferroic Metallic Monolayer Cu(CrSe2)2

### v1

@@ Abstract
The two-dimensional (2D) Cu(CrSe2)2 monolayer stands out for its combined ferromagnetic (FM), ferroelectric (FE), and metallic properties, marking itself as a prominent 2D multiferroic metal. This work studies those properties and the relevant physics, using density functional calculations, Monte Carlo simulations, and molecular dynamics. Our results show that Cu(CrSe2)2 monolayer is in the Cr3+ state with = 3/2 and Cu1+ with = 0. A ligand hole in the Se 4 orbitals gives rise to metallic behavior and enhances the FM coupling between the local Cr3+ = 3/2 spins. The observed in-plane magnetic anisotropy primarily arises from exchange anisotropy, which is associated with the Cr-Se-Cr itinerant ferromagnetism. In contrast, both single-ion anisotropy and shape magnetic anisotropy contribute negligibly. The Dzyaloshinskii-Moriya interaction is also quite weak, only about 3% of the intralayer exchange parameters. Our Monte Carlo simulations show a FM Curie temperature () of 190 K. Moreover, the monolayer exhibits a vertical FE polarization of 1.79 pC/m and a FE polarization switching barrier of 182 meV/f.u., and the FE state remains stable above 800 K as shown by molecular dynamics simulations. Furthermore, a magnetoelectric coupling is partially manifested by a magnetization rotation from in-plane to out-of-plane associated with a FE-to-paraelectric transition. The magnetization rotation can also be induced by either hole or electron doping, and the hole doping increases the up to 238 K. In addition, tensile strain reduces the FE polarization but enhances to 290 K, while a compressive strain gives an opposite effect. Therefore, the multiferroic metallic Cu(CrSe2)2 monolayer may be explored for advanced multifunctional electronic devices.
Introduction
@@
Research on two-dimensional (2D) materials has substantially expanded due to the discovery of 2D ferromagnetic (FM) materials, including CrI3 [1] and Cr2Ge2Te6 [2], and 2D ferroelectric (FE) materials, such as CuInP2S6 [3] and SnTe [4]. The Mermin-Wagner theorem posits that long-range magnetic order cannot be sustained at any finite temperature in 2D isotropic Heisenberg spin systems [5]. This limitation of 2D systems challenges the conventional understanding of magnetic behavior and underscores the importance of magnetic anisotropy (MA) in stabilizing their magnetic states [6, 7]. Ferroelectricity, a characteristic of certain insulators, displays spontaneous electric polarization reversible by an external electric field [8, 9]. Anderson and Blount introduced the concept of “FE metal” [10]. During a symmetry-breaking structural transition, it is hypothesized that a metal could generate FE. This hypothesis has been recently confirmed through 3D LiOsO3 and NdNiO3 [11, 12]. In the realm of 2D structures, FE metals are exceedingly rare [13]. To date, the only evidence is the demonstrated ability to reverse polarization in a 2D WTe2 FE metal by applying an electric field [14]. When magnetization and FE metallicity are combined in the 2D limit, the result is a “2D multiferroic metal”. This material not only exhibits characteristics typical of bulk multiferroics, such as phase transitions [15], magneto-electric coupling [16], and unconventional optical responses [17], but also offers unique advantages due to its monolayer thickness. These properties can be effectively manipulated using magnetic and electric fields, doping, Moir engineering, or the creation of heterostructures [18, 7, 19, 20, 21, 22].
@@
Very recently, the 2D Cu(CrSe2)2 monolayer was derived from CuCrSe2 bulk [23, 24] using a redox-controlled electrochemical exfoliation method [25, 26]. As depicted in Fig. 1, this monolayer features a Cu atom layer sandwiched between two CrSe2 layers, where Cr-cations form triangular nets through edge-shared CrSe6 octahedra. Each Cu atom forms a tetrahedral bond, connecting to one Se atom from the top CrSe2 layer and three Se atoms from the bottom layer. Experimental investigations have confirmed that the Cu(CrSe2)2 monolayer exhibits FM behavior with a Curie temperature () of 125 K and notable in-plane magnetization [25]. Meanwhile, the FE phase transition temperature for ultrathin CuCrSe2 has been measured at approximately 800 K [27]. Previously, the multiferroic behavior was also theoretically predicted to be above 300 K [28], which is in contrast with the experimental of 125 K. More recently, a study has reported that single-layer Cu(CrSe2)2 not only maintains ferroelectricity above room temperature but also displays ferromagnetism at 120 K [26], further underscoring its intriguing multiferroic nature.
@@
Here, we are motivated to study those multiferroic properties of Cu(CrSe2)2 monolayer and particularly the relevant physics, using density functional calculations, Monte Carlo simulations, and molecular dynamics. Our results show that the itinerant Se holes enhance the Cr-Cr FM coupling and the exchange anisotropy determines the observed in-plane MA. Our MC simulations give the FM of 190 K, which is moderately higher than the experimental 120-125 K [25, 26]. Moreover, the monolayer exhibits a vertical FE polarization of 1.79 pC/m and a FE switching barrier of 182 meV/f.u., and the FE state remains stable above 800 K as proven by our molecular dynamics simulations. Furthermore, we predict that the magnetization would rotate from in-plane to out-of-plane upon a FE-to-paraelectric transition, showing a magnetoelectric coupling. The magnetization rotation can also be induced by either hole or electron doping, and a hole doping would increase up to 238 K. A tensile strain would reduce the FE polarization but elevate up to 290 K. Thus, our work provides a systematic and comprehensive investigation of the multiferroic properties of Cu(CrSe2)2 monolayer, potentially aiding in the exploration of its multifunctionality in spintronics.
@@
We perform density functional theory (DFT) calculations using the Vienna ab simulation package (VASP) [29]. The kinetic energy cutoff for the plane wave expansion is set to 500 eV. The -centered mesh of 1515 is used for the 11 unit cell and (158) k-mesh for the 12 supercell. The setup of kinetic energy and k-point sampling are carefully tested to ensure the convergence of the MA results within a few percent, as shown in Table S1 in supporting information (SI). Atomic positions are fully relaxed until the force on each atom is less than 0.01 eV/Å , and the total energy minimization is performed with a tolerance of 10-7 eV. The DFT-optimized lattice constants = = 3.63 Å with space group are used for the Cu(CrSe2)2 monolayer, which are almost the same as the experimental bulk values of = = 3.69 Å [23]. A vacuum space larger than 15 Å is used.
To account for the electron correlation of the Cu (Cr) states, the generalized gradient approximation plus Hubbard (GGA + ) [30] calculations are performed with Hubbard = 6.0 eV (4.0 eV) and Hund exchange = 0.9 eV (0.9 eV). The SOC is included for Cu, Cr, and Se atoms by the second-variation method with scalar relativistic wave functions. We also tested eV and eV (see Table S2 in the SI). Although increasing results in some quantitative changes in the exchange parameters and the magnetic anisotropy energy (MAE), our conclusions remain unaffected, , the estimated of the Cu(CrSe2)2 monolayer undergoes only a small change from 190 K to 178 K. Dipole moment correction is applied to evaluate the vertical polarization and the FE switching path is calculated using the climbing image nudged elastic band (NEB) method [31]. The phonon spectrum is calculated using Phonopy software interfaced with VASP [32]. In addition, molecular dynamics (AIMD) simulations are performed to study the stability of the FE polarization in a 22 supercell over 5 ps at different temperatures. The magnetic phase transition of the Cu(CrSe2)2 monolayer is studied using Monte Carlo (MC) simulations. In our MC simulations, a lattice is chosen. During the simulation step, each spin is randomly rotated in three-dimensional space. The spin dynamical process is studied by the classical Metropolis methods [33].
@@
.1 Electronic states and magnetic properties
@@
With the confirmed easy in-plane magnetization, we now turn to the estimation of the magnetic interactions. The Cu(CrSe2)2 monolayer is characterized by Cr3+ = 3/2 and Cu1+ charge states, with a metallic ligand hole predominantly on the Se 4 band. One thus may expect an itinerant FM ground state. To substantiate this, we employed GGA+SOC+ calculations on a 12 supercell, examining three additional antiferromagnetic (AF) states with in-plane magnetization, as depicted in Fig. S1 in SI. Indeed, the FM metallic state is more stable than the three AF states, as seen in Table S5 in SI. Our results indicate that both CrSe2 layers and the Cu-connected interlayer exhibit FM coupling. In order to obtain the exchange parameters, we introduced three specific ones: for the top layer of CrSe2, for the bottom layer, and for the interactions between the CrSe2 layers, as illustrated in Fig. S1 in the SI. Considering that the nearest-neighbor Cr ions in the CrSe2 layers are six-coordinated and those in the interlayer are three-coordinated, we express as follows the magnetic exchange energies for the FM ground state and three AF states per formula unit, with an assumption of for a FM pair of Cr3+ = 3/2 ions:
@@
The Cu(CrSe2)2 monolayer, with the space group 31 and lacking spatial inversion symmetry, exhibits also Dzyaloshinskii-Moriya (DM) interactions [39, 40]. We now estimate the DM interactions using the four-state energy mapping approach [41]. As an example, we calculate the interactions between a pair of nearest Cr ions along the -axis. The results for the DM interaction are [0.09, –0.08, 0.09] meV, with = 0.15 meV for the top layer and [0.17, –0.11, 0.00] meV, with = 0.20 meV for the bottom layer, see Tables S3 and S4 in SI. The / ratios on the top and bottom Cr layers are 2.6% and 3.0%, respectively. The DM interactions are much weaker than the intralayer exchange parameters in this material and can therefore be neglected in this work.
@@
The first term describes the above intralayer and interlayer isotropic Heisenberg exchange, and the second term represents the MAE with = 0.04 meV for the easy in-plane magnetization. Our MC simulations show that the Cu(CrSe2)2 monolayer has a of 190 K, which is much lower than the previous theoretical prediction of over 300 K [28] but is a moderate overestimation, compared with the experimental values of 120-125 K [25, 26].
Considering the FM metallic nature of the Cu(CrSe2)2 monolayer and possible long-range magnetic couplings, we also calculate the exchange parameters up to the third-nearest neighbors (), as detailed in Figure S2, Equation S1, and Table S6 in the SI. These results show that although the specific exchange parameters vary upon the inclusion of the different magnetic couplings in the spin Hamiltonian, here the and for both the top and bottom CrSe2 layers are nearly one order of magnitude weaker than the above and , and a core result–the estimated remains almost unchanged, being 185 K given by our MC simulations.
.2 Ferroelectric properties
In addition to its 2D FM metallic behavior, the Cu(CrSe2)2 monolayer may also exhibit FE properties due to its space group and lack of spatial inversion symmetry. As seen in Fig. 1(a), the originally centrosymmetric bilayer of CrSe2 is disrupted by the intercalation of Cu ions, resulting in an out-of-plane FE polarization [42]. Our calculations indicate that the monolayer exhibits an FE polarization of approximately 1.79 pC/m in the negative direction along the c-axis, representing a vertical downward FE polarization (P). In 2D metallic systems, electrons move freely within the basal plane but face constraints in the vertical direction. This characteristic might enable the persistence of out-of-plane polarization, which could be altered by a perpendicular electric field, as suggested for the 2D FE metal WTe2 [14].
@@
To explore the stability of ferroelectricity under finite temperatures and to ascertain its FE phase transition temperature, we conducted AIMD simulations. The results indicate that the probability of FE polarization exhibits a singular peak and adheres to the normal (Gaussian) distribution at each temperature, represented as (,). Herein, represents the mean value, and denotes the standard deviation value. As seen in Fig. 3, the material exhibits robust ferroelectricity at 300 K and maintains substantial ferroelectricity at 800 K, evolving into a paraelectric (PE) state at 900 K. This observation aligns well with the experimental result that the measured ferroelectric phase transition temperature of ultrathin CuCrSe2 is about 800 K [27]. Hence, Cu(CrSe2)2 monolayer could be an optimal 2D room-temperature FE material, in addition to its FM metallic behavior, and it would then be a new 2D multiferroic metal.
@@
The reversal barrier of FE polarization is crucial for its practical applications. Here, we explore two possible reversal pathways. In the Cu(CrSe2)2 structure, shifting the Cu ions can flip the state from a P polarization to an energetically equivalent P polarization, as illustrated in Fig. 4. To elucidate the specific pathway for this switching in the Cu(CrSe2)2 monolayer, we utilized the NEB method [31]. The results, depicted in Fig. 4(a), reveal that during the transition phase, Cu ions adopt a two-coordinated state, representing the paraelectric phase (PE′) as shown in Fig. 4(a). With a calculated switching barrier of 328 meV for each Cu ion, this large reversal barrier suggests that a high electric field would likely be required to reverse the polarization. Instead, the second reversal pathway is explored and turns out to be much lower and could then be more economic for an electric switch. Here, the CrSe2 layers exhibit in-plane shifts as indicated by the pink arrows, and each Cu ion is coordinated with six Se atoms to form the CuSe6 octahedron as shown in Fig. 4(b). Our calculations indicate a much reduced switching barrier of 182 meV/f.u., well comparable with those of conventional 3D FE materials such as PbTiO3 (200 meV/f.u.) [43]. It is worth noting that this intermediate PE state is a local energy minimum, with its dynamic stability being corroborated also by phonon calculations, as seen in Fig S3 in SI.
.3 Magnetoelectric, doping, and strain effects
@@
Our above results suggest that the Cu(CrSe2)2 monolayer is not only a 2D FM metal but also FE, and thus it is added as a new case into the rare 2D multiferroic metals. Next, we examine its potential magnetoelectric effects during the FE-PE transition. Our results show that the MA’s direction switches from the in-plane to out-of-plane during the FE-PE transition, as seen in Fig. 5. Moreover, the exhibits variations, peaking around the PE phase. To understand the switch of MA axes, we computed the MAE contribution from each atomic layer. As observed from Figs. 5(b) and (c), the primary source of MAE in the FE state lies with the Se ion layers, being consistent with our above results. However, we observed significant changes in the MAE across the Se ion layers during the FE-PE transition. For example, the MAE of the second Se layer changes from the in-plane one of 1.12 meV/f.u. to the out-of-plane one of 0.31 meV/f.u. It is this change that plays a crucial role in the MA switch from in-plane to out-of-plane associated with the FE-PE transition.
@@
Considering the material’s FM itinerant nature and the manipulability of 2D materials, we further explore the control of its multiferroic properties through carrier dopings [44, 45]. As illustrated in Figs. 1(c), a finite hole doping would increase the density of band states at the Fermi level and thus favors the itinerant FM, whereas a partial electron doping could have an opposite effect and may even drive the system into a semiconductor. In our calculations, the hole (electron) doping is simulated using the virtual crystal approximation (VCA) [46]. As expected, hole doping enhances FM coupling, raising from 190 K to 238 K with 0.8 holes per formula unit, as shown in Fig. 6. Notably, with the hole doping, the easy MA direction changes from the otherwise in-plane to the out-of-plane.
@@
We investigated the magnetoelectric properties and the relevant physics in the new 2D multiferroic metallic Cu(CrSe2)2 monolayer, using density functional calculations, Monte Carlo simulations, and molecular dynamics. Our results show that the monolayer is in the Cr3+ = 3/2 and Cu+ = 0 state. The ligand Se 4 holes enhance the FM metallicity with the about 190 K, and the Se -Cr exchange determines the in-plane magnetic anisotropy (MA), in which the single ion anisotropy and shape MA are all much weaker or even negligible. Moreover, we predict that this monolayer retains the vertical FE polarization above 800 K. Therefore, Cu(CrSe2)2 monolayer seems to be a quite stable multiferroic metal. Furthermore, our results predict that the magnetoelectric properties could be tuned by electric field, carrier doping, and lattice strain: a MA switch from the in-plane to out-of-plane upon the FE-PE transition or hole doping, and hole doping or tensile strain enhanced even up to room temperature. Thus, Cu(CrSe2)2 monolayer is a new and rare 2D multiferroic metal and could have promising spintronic applications.

### v2

@@ Abstract
The two-dimensional (2D) Cu(CrSe2)2 monolayer stands out for its combined ferromagnetic (FM), ferroelectric (FE), and metallic properties, marking itself as a prominent 2D multiferroic metal. This work studies those properties and the relevant physics, using density functional calculations, Monte Carlo simulations, and molecular dynamics. Our results show that Cu(CrSe2)2 monolayer is in the Cr3+ state with = 3/2 and Cu1+ with = 0. A ligand hole in the Se 4 orbitals gives rise to metallic behavior and enhances the FM coupling between the local Cr3+ = 3/2 spins. The observed in-plane magnetic anisotropy primarily arises from exchange anisotropy, which is associated with the Cr-Se-Cr itinerant ferromagnetism. In contrast, both single-ion anisotropy and shape magnetic anisotropy contribute negligibly. The Dzyaloshinskii-Moriya interaction is also quite weak, only about 3% of the intralayer exchange parameters. Our Monte Carlo simulations show a FM Curie temperature () of 190 K. Moreover, the monolayer exhibits a vertical FE polarization of 1.79 pC/m and a FE polarization switching barrier of 182 meV/f.u., and the FE state remains stable above room temperature as shown by molecular dynamics simulations. Furthermore, a magnetoelectric coupling is partially manifested by a magnetization rotation from in-plane to out-of-plane associated with a FE-to-paraelectric transition. The magnetization rotation can also be induced by either hole or electron doping, and the hole doping increases the up to 238 K. In addition, tensile strain reduces the FE polarization but enhances to 290 K, while a compressive strain gives an opposite effect. Therefore, the multiferroic metallic Cu(CrSe2)2 monolayer may be explored for advanced multifunctional electronic devices.
@@
Research on two-dimensional (2D) materials has substantially expanded due to the discovery of 2D ferromagnetic (FM) materials, including CrI3 [1] and Cr2Ge2Te6 [2], and 2D ferroelectric (FE) materials, such as CuInP2S6 [3] and SnTe [4]. The Mermin-Wagner theorem posits that long-range magnetic order cannot be sustained at any finite temperature in 2D isotropic Heisenberg spin systems [5]. This limitation of 2D systems challenges the conventional understanding of magnetic behavior and underscores the importance of magnetic anisotropy (MA) in stabilizing their magnetic states [6, 7]. Ferroelectricity, a characteristic of certain insulators, displays spontaneous electric polarization reversible by an external electric field [8, 9]. Anderson and Blount introduced the concept of “FE metal” [10]. During a symmetry-breaking structural transition, it is hypothesized that a metal could generate FE. This hypothesis has been recently confirmed through 3D LiOsO3 and NdNiO3 [11, 12]. In the realm of 2D structures, FE metals are exceedingly rare [13]. To date, the only evidence is the demonstrated ability to reverse polarization in a 2D WTe2 FE metal by applying an electric field [14]. When magnetization and FE metallicity are combined in the 2D limit, the result is a “2D multiferroic metal”. This material not only exhibits characteristics typical of bulk multiferroics, such as phase transitions [15], magneto-electric coupling [16], and unconventional optical responses [17], but also offers unique advantages due to its monolayer thickness. These properties can be effectively manipulated using magnetic and electric fields, doping, Moir engineering, or the creation of heterostructures [7, 18, 19, 20, 21, 22, 23, 24].
@@
Very recently, the 2D Cu(CrSe2)2 monolayer was derived from CuCrSe2 bulk [25, 26] using a redox-controlled electrochemical exfoliation method [27, 28]. As depicted in Figure 1, this monolayer features a Cu atom layer sandwiched between two CrSe2 layers, where Cr-cations form triangular nets through edge-shared CrSe6 octahedra. Each Cu atom forms a tetrahedral bond, connecting to one Se atom from the top CrSe2 layer and three Se atoms from the bottom layer. Experimental investigations have confirmed that the Cu(CrSe2)2 monolayer exhibits FM behavior with a Curie temperature () of 125 K and notable in-plane magnetization [27]. Meanwhile, the FE phase transition temperature for ultrathin CuCrSe2 has been measured at approximately 800 K [29]. Previously, the multiferroic behavior was also theoretically predicted to be above 300 K [30], which is in contrast with the experimental of 125 K. More recently, a study has reported that single-layer Cu(CrSe2)2 not only maintains ferroelectricity above room temperature but also displays ferromagnetism at 120 K [28], further underscoring its intriguing multiferroic nature.
@@
Here, we are motivated to study those multiferroic properties of Cu(CrSe2)2 monolayer and particularly the relevant physics, using density functional calculations, Monte Carlo (MC) simulations, and molecular dynamics. Our results show that the itinerant Se holes enhance the Cr-Cr FM coupling and the exchange anisotropy determines the observed in-plane MA. Our MC simulations give the FM of 190 K, which is moderately higher than the experimental 120-125 K [27, 28]. Moreover, the monolayer exhibits a vertical FE polarization of 1.79 pC/m and a FE switching barrier of 182 meV/f.u., and the FE state remains stable above room temperature as proven by our molecular dynamics simulations. Furthermore, we predict that the magnetization would rotate from in-plane to out-of-plane upon a FE-to-paraelectric transition, showing a magnetoelectric coupling. The magnetization rotation can also be induced by either hole or electron doping, and a hole doping would increase up to 238 K. A tensile strain would reduce the FE polarization but elevate up to 290 K. Thus, our work provides a systematic and comprehensive investigation of the multiferroic properties of Cu(CrSe2)2 monolayer, potentially aiding in the exploration of its multifunctionality in spintronics.
@@
We perform density functional theory (DFT) calculations using the Vienna ab simulation package (VASP) [31]. The kinetic energy cutoff for the plane wave expansion is set to 500 eV. The -centered mesh of 1515 is used for the 11 unit cell and (158) k-mesh for the 12 supercell. The setup of kinetic energy and k-point sampling are carefully tested to ensure the convergence of the MA results within a few percent, as shown in Table S1 in supporting information (SI). Atomic positions are fully relaxed until the force on each atom is less than 0.01 eV/Å , and the total energy minimization is performed with a tolerance of 10-7 eV. The DFT-optimized lattice constants = = 3.63 with space group are used for the Cu(CrSe2)2 monolayer, which are almost the same as the experimental bulk values of = = 3.69 [25]. A vacuum space larger than 15 Å is used.
To account for the electron correlation of the Cu (Cr) states, the generalized gradient approximation plus Hubbard (GGA + ) [32] calculations are performed with Hubbard = 6.0 eV (4.0 eV) and Hund exchange = 0.9 eV (0.9 eV). The Hubbard was also used in the previous calculations of the FM metal CrO[33]. The SOC is included for Cu, Cr, and Se atoms by the second-variation method with scalar relativistic wave functions. We also tested eV and eV (see Table S2 in the SI). Although increasing results in some quantitative changes in the exchange parameters and the magnetic anisotropy energy (MAE), our conclusions remain unaffected, , the estimated of the Cu(CrSe2)2 monolayer undergoes only a small change from 190 K to 178 K. Note that if the Hubbard is not used, the plain GGA calculations (see Fig. S1) overestimate the itineracy of the Cr electrons and give the enhanced FM exchange and exchange anisotropy (see Table S2), and then the MC simulation gives of 360 K, being much higher than the experimental 120-125 K [27, 28]. Dipole moment correction is applied to evaluate the vertical polarization and the FE switching path is calculated using the climbing image nudged elastic band (NEB) method [34]. The phonon spectrum is calculated using Phonopy software interfaced with VASP [35]. In addition, molecular dynamics (AIMD) simulations are performed to study the stability of the FE polarization in a 44 supercell over 30 ps at different temperatures. The magnetic phase transition of the Cu(CrSe2)2 monolayer is studied using MC simulations. In our MC simulations, a lattice is chosen. During the simulation step, each spin is randomly rotated in three-dimensional space. The spin dynamical process is studied by the classical Metropolis methods [36].
@@
Electronic states and magnetic properties
@@
With the confirmed easy in-plane magnetization, we now turn to the estimation of the magnetic interactions. The Cu(CrSe2)2 monolayer is characterized by Cr3+ = 3/2 and Cu1+ charge states, with a metallic ligand hole predominantly on the Se 4 band. One thus may expect an itinerant FM ground state. To substantiate this, we employed GGA+SOC+ calculations on a 12 supercell, examining three additional antiferromagnetic (AF) states with in-plane magnetization, as depicted in Fig. S2 in SI. Indeed, the FM metallic state is more stable than the three AF states, as seen in Table S3 in SI. Our results indicate that both CrSe2 layers and the Cu-connected interlayer exhibit FM coupling. In order to obtain the exchange parameters, we introduced three specific ones: for the top layer of CrSe2, for the bottom layer, and for the interactions between the CrSe2 layers, as illustrated in Fig. S2 in SI. Considering that the nearest-neighbor Cr ions in the CrSe2 layers are six-coordinated and those in the interlayer are three-coordinated, we express as follows the magnetic exchange energies for the FM ground state and three AF states per formula unit, with an assumption of for a FM pair of Cr3+ = 3/2 ions:
@@
The Cu(CrSe2)2 monolayer, with the space group 31 and lacking spatial inversion symmetry, exhibits also Dzyaloshinskii-Moriya (DM) interactions [42, 43]. We now estimate the DM interactions using the four-state energy mapping approach [44]. As an example, we calculate the interactions between a pair of nearest Cr ions along the -axis. The results for the DM interaction are [0.09, –0.08, 0.09] meV, with = 0.15 meV for the top layer and [0.17, –0.11, 0.00] meV, with = 0.20 meV for the bottom layer, see Tables S4 and S5 in SI. The / ratios on the top and bottom Cr layers are 2.6% and 3.0%, respectively. The DM interactions are much weaker than the intralayer exchange parameters in this material and can therefore be neglected in this work.
@@
The sum over runs over the entire lattice of Cr atoms, and runs over the six nearest Cr neighbors of each Cr atom. The first term describes the intralayer and interlayer isotropic Heisenberg exchange interaction. The second term represents the exchange anisotropy, which is caused by the SOC effect of the ligand Se holes mediating the FM Cr-Cr coupling, giving the energy difference between in-plane and out-of-plane magnetizations. In our calculations, the exchange anisotropy constant is determined to be 0.04 meV, and it corresponds to the easy in-plane magnetization. Our MC simulations show that the Cu(CrSe2)2 monolayer has a of 190 K, which is much lower than the previous theoretical prediction of over 300 K [30] but is a moderate overestimation, compared with the experimental values of 120-125 K [27, 28].
Considering the FM metallic nature of the Cu(CrSe2)2 monolayer and possible long-range magnetic couplings, we also calculate the exchange parameters up to the third-nearest neighbors (), as detailed in Fig. S3, Equation S1, and Table S6 in SI. These results show that although the specific exchange parameters vary upon the inclusion of the different magnetic couplings in the spin Hamiltonian, here the and for both the top and bottom CrSe2 layers are nearly one order of magnitude weaker than the above and , and a core result–the estimated remains almost unchanged, being 185 K given by our MC simulations.
Ferroelectric properties
In addition to its 2D FM metallic behavior, the Cu(CrSe2)2 monolayer may also exhibit FE properties due to its space group and lack of spatial inversion symmetry. As seen in Fig. 1(a), the originally centrosymmetric bilayer of CrSe2 is disrupted by the intercalation of Cu ions, resulting in an out-of-plane FE polarization [45]. In this study, the dipole moment is computed directly by integrating the charge density (including both electronic and ionic contributions) of the metallic thin film. The FE polarization is calculated by applying a dipole moment correction to evaluate the vertical polarization of the Cu(CrSe2)2 monolayer. This method has already been justified in the study of the similar FE metallic WTe2 bilayer [46]. Our calculations indicate that the monolayer exhibits an FE polarization of approximately 1.79 pC/m in the negative direction along the c-axis, representing a vertical downward FE polarization (P). In 2D metallic systems, electrons move freely within the basal plane but face constraints in the vertical direction. This characteristic might enable the persistence of out-of-plane polarization, which could be altered by a perpendicular electric field, as suggested for the 2D FE metal WTe2 [14].
@@
To explore the stability of ferroelectricity under finite temperatures and to ascertain its FE phase transition temperature, we conducted AIMD simulations. We statistically analyze the thermal equilibrium states between 5 ps and 30 ps, as shown in Fig. S5, to calculate the probability of FE polarization. The results indicate that the probability of FE polarization exhibits a singular peak and adheres to the normal (Gaussian) distribution at each temperature, represented as (,). Herein, represents the mean value, and denotes the standard deviation value. As seen in Fig. 3, the material exhibits robust ferroelectricity above room temperature (even about 500 K). This result agrees well with the most recent experimental observation of the FE phase in monolayer CuCrSe2 above room temperature [28]. Hence, Cu(CrSe2)2 monolayer could be an optimal 2D room-temperature FE material, in addition to its FM metallic behavior, and it would then be a new 2D multiferroic metal.
@@
The reversal barrier of FE polarization is crucial for its practical applications. Here, we explore three possible reversal pathways. In the Cu(CrSe2)2 structure, shifting the Cu ions can flip the state from a P polarization to an energetically equivalent P polarization, as illustrated in Fig. 4. To elucidate the specific pathway for this switching in the Cu(CrSe2)2 monolayer, we utilized the NEB method [34]. The results, depicted in Fig. 4(a), reveal that during the transition phase, Cu ions adopt a two-coordinated state, representing the paraelectric phase (PE′) as shown in Fig. 4(a). With a calculated switching barrier of 328 meV for each Cu ion, this large reversal barrier suggests that a high electric field would likely be required to reverse the polarization. Second, as a switching pathway via the anti-FE structure is possible [47], we calculate it and find the switching barrier is 203 meV/f.u. Furthermore, the third pathway is explored and turns out to be most economic for an electric switch. Here, the CrSe2 layers exhibit in-plane shifts as indicated by the pink arrows, and each Cu ion is coordinated with six Se atoms to form the CuSe6 octahedron as shown in Fig. 4(b). Our calculations indicate a much reduced switching barrier of 182 meV/f.u., well comparable with those of conventional 3D FE materials such as PbTiO3 (200 meV/f.u.) [48]. It is worth noting that this intermediate PE state is a local energy minimum, with its dynamic stability being corroborated also by phonon calculations, as seen in Fig. S4 in SI.
Magnetoelectric, doping, and strain effects
@@
Our above results suggest that the Cu(CrSe2)2 monolayer is not only a 2D FM metal but also FE, and thus it is added as a new case into the rare 2D multiferroic metals. As seen in Fig. 4(b), although the PE phase has a higher energy than the FE phase, it is a local energy minimum. Thus, the PE state could stably exist during the FE switching. Next, we examine a potential magnetoelectric effect, namely, varying magnetism associated with the FE-PE transition driven by an electric field. Our results show that the MA’s direction switches from the in-plane to out-of-plane during the FE-PE transition, as seen in Fig. 5. Moreover, the exhibits variations, peaking around the PE phase. To understand the switch of MA axes, we computed the MAE contribution from each atomic layer. As observed from Figs. 5(b) and (c), the primary source of MAE in the FE state lies with the Se ion layers, being consistent with our above results. However, we observed significant changes in the MAE across the Se ion layers during the FE-PE transition. For example, the MAE of the second Se layer changes from the in-plane one of 1.12 meV/f.u. to the out-of-plane one of 0.31 meV/f.u. It is this change that plays a crucial role in the MA switch from in-plane to out-of-plane associated with the FE-PE transition.
@@
Considering the material’s FM itinerant nature and the manipulability of 2D materials, we further explore the control of its multiferroic properties through carrier dopings [49, 50]. As illustrated in Figs. 1(c), a finite hole doping would increase the density of band states at the Fermi level and thus favors the itinerant FM, whereas a partial electron doping could have an opposite effect. In our calculations, the hole (electron) doping is simulated using the virtual crystal approximation (VCA) [51]. As expected, hole doping enhances FM coupling, raising from 190 K to 238 K with 0.8 holes per formula unit, as shown in Fig. 6. Notably, with the hole doping, the easy MA direction changes from the otherwise in-plane to the out-of-plane.
@@
We investigated the magnetoelectric properties and the relevant physics in the new 2D multiferroic metallic Cu(CrSe2)2 monolayer, using density functional calculations, Monte Carlo simulations, and molecular dynamics. Our results show that the monolayer is in the Cr3+ = 3/2 and Cu+ = 0 state. The ligand Se 4 holes enhance the FM metallicity with the about 190 K, and the Se -Cr exchange determines the in-plane MA, in which the single ion anisotropy and shape MA are all much weaker or even negligible. Moreover, we predict that this monolayer retains the vertical FE polarization above room temperature. Therefore, Cu(CrSe2)2 monolayer seems to be a quite stable multiferroic metal. Furthermore, our results predict that the magnetoelectric properties could be tuned by electric field, carrier doping, and lattice strain: a MA switch from the in-plane to out-of-plane upon the FE-PE transition or hole doping, and hole doping or tensile strain enhanced even up to room temperature. Thus, Cu(CrSe2)2 monolayer is a new and rare 2D multiferroic metal and could have promising spintronic applications.